OpenStudy (anonymous):

write the given expression in terms of x and y only tan(sin−1 x + cos−1 y)

OpenStudy (souvik):

let sin\(\theta\)=x and cos\(\alpha\)=y now the expression is tan(\(\theta+\alpha\))

OpenStudy (anonymous):

and then? what is the value of sin? cos?

OpenStudy (souvik):

sin\(\theta\)=x so cos\(\theta\)=\(\sqrt{1-x^2}\) now tan(\(\theta+\alpha\)=tan\(\theta\)+tan\(\alpha\)/(1-tan\(\theta\)tan\(\alpha\))

OpenStudy (anonymous):

tangent will be x/sqrt of 1-x^2?

OpenStudy (souvik):

yes...good

OpenStudy (anonymous):

how come that cos= sqrt of 1-x^2?

OpenStudy (souvik):

do u know this? sin^2\(\theta\)+cos^2\(\theta\)=1

OpenStudy (anonymous):

yes

OpenStudy (souvik):

i get it from there...

OpenStudy (anonymous):

but sine function has -1 exponent.

OpenStudy (souvik):

sin\(\theta\)=x sin^2\(\theta\)=x^2 sin^2θ+cos^2θ=1 or x^2+cos^2\(\theta\)=1 or cos\(\theta\)=\(\sqrt{1-x^2}\)

OpenStudy (anonymous):

pls help i cant understand. the -1 exponent? I will disregard it?

OpenStudy (souvik):

if sin\(\theta\)=x then \(\theta\)=sin^-1x

OpenStudy (anonymous):

what will be the equivalent of tan(\[\theta\]

OpenStudy (anonymous):

pls help show me the answer

OpenStudy (souvik):

lets do it together...

OpenStudy (souvik):

cosα=y sinα=?

OpenStudy (anonymous):

sin\[\sin \alpha =\sqrt{1-y ^{2}}\]

OpenStudy (anonymous):

? am I right?

OpenStudy (souvik):

now tan \(\alpha\)=?

OpenStudy (anonymous):

\[\frac{ y }{ \sqrt{} }\]

OpenStudy (anonymous):

\[\frac{ \sqrt{1-y ^{2}}}{y }\]

OpenStudy (souvik):

now tan(θ+α)=tanθ+tanα/(1-tanθtanα) just put what u get fo tan\(\alpha\) and tan\(\theta\)

OpenStudy (anonymous):

I will disregard the -1 exponent of sine?

OpenStudy (souvik):

tanθ+tanα/(1-tanθtanα) put the tan\(\theta\) and tan\(\alpha\) here

OpenStudy (anonymous):

then after i put the given? whats next? simplify?

OpenStudy (souvik):

then simplify